### v1

@@ 3.2 Results
pinnA is a fully-connected neural network trained along the time axis using a time-marching approach. The configuration consists of 8 hidden layers with 128 neurons per layer. The global time interval is partitioned into 20 consecutive windows of width $\Delta t_{\text{window}}=0.10$, and each window is trained independently. The loss function is defined as a weighted sum of the PDE, boundary, initial condition, window-interface continuity, and Poynting-vector-based energy conservation terms.
The full configuration of the model is summarized in the table below.
Note WINDOW_SIZE = 0.10
BETA = 1.0 # Aspect ratio of cavity
A = 1.0 # Gaussian pulse amplitude
SIGMA = 0.1 # Gaussian pulse width
X_MIN, X_MAX = 0.0, 1.0
Y_MIN, Y_MAX = 0.0, BETA
T_MIN, T_MAX = 0.0, 2.0
# Sampling points (per window)
N_COLLOCATION = int(50000 / BETA) // len(TIME_WINDOWS) * 2
N_BOUNDARY = int(10000 / BETA) // len(TIME_WINDOWS) * 2
N_INITIAL = int(3000 / BETA)
W_PDE = 1.0
W_BC = 50.0 / BETA
W_IC = 100.0
W_INTERFACE = 75.0
W_POYNTING = 50.0
N_HIDDEN_LAYERS = 8
N_NEURONS = 128
N_ADAM_EPOCHS = 1500 (per window)
N_LBFGS_EPOCHS = 300 (per window)
LEARNING_RATE_ADAM = 1e-3
LEARNING_RATE_LBFGS = 0.5
@@
Field-level accuracy of pinnA relative to FDTD is assessed using the Normalized Root Mean Square Error (NRMSE) and the relative $L^{2}$ norms defined in the previous section. The results for selected time steps are reported in Table 1. The reported average values are computed not only over the time steps shown in the table, but as an average over all time steps used in the analysis, $t=[0.00,0.05,0.10,0.15,\ldots,2.00]$.
As shown in the table, the NRMSE values for all three field components remain on the order of 0.1–0.2% throughout the full time interval. Even the relative $L^{2}$ error reaches only about $1.33\%$ in the worst case and stays around $1.01\%$ on average.
@@
Since the values in the table and the figure are computed on a collocated grid for the PINN solution, FDTD and PINN are additionally evaluated on a common Yee grid to confirm consistency. Under the staggered-grid layout, the mean absolute difference over time is $3.74\times 10^{-6}$, and the maximum difference is $9.07\times 10^{-6}$. When normalized by the reference energy, the mean relative mismatch is only $0.024\%$, and the maximum mismatch is $0.057\%$. In other words, the discrete energy functional of the PINN tracks the discrete FDTD energy across the entire spatio-temporal domain at a level well below $10^{-3}$ relative error.
@@
Resolving the Cumulative Energy Drift
Two characteristic issues of time-marched PINNs, cumulative energy drift and discontinuities at window transitions, are suppressed in pinnA through two complementary mechanisms. The purpose of this section is to briefly describe the existence of these mechanisms and their functional role; a detailed comparison of the Poynting variants is provided in the ablation study.
a) Poynting Loss for cumulative energy drift: In pinnA, energy conservation is reinforced by adding the Poynting-based constraint defined in section 2 to the loss function as a physics-based regularizer. In the final configuration, the Poynting constraint is applied in a local form, as an additional physics constraint evaluated at the same collocation points as the Maxwell PDE residual. This design yields a regularization mechanism that directly influences the solution dynamics within each window, rather than leaving energy behavior to a single global integral check. The differences between global and local Poynting formulations, and the effect of this choice on energy drift, are presented in detail in section 3.3 together with the corresponding ablation experiments.
b) Interface Continuity Loss for discontinuities at window transitions: To reliably stitch successive windows, pinnA employs a two-stage continuity mechanism. The first stage is the IC continuity loss applied at the start of each window. In the first window, the initial-condition term enforces the physical initial state by prescribing a Gaussian pulse and zero magnetic field. In subsequent windows, the same term no longer targets an analytic profile, but instead matches the trained PINN solution from the previous window. Accordingly, the $E_{z}$, $H_{x}$, and $H_{y}$ fields at the beginning of window $k$ are penalized in terms of squared error against the fields at the terminal time of window k1k-1. The fields imported from the previous window are ‘detached‘ during backpropagation so that they serve as fixed targets; gradients therefore flow only through the parameters of the active window, keeping the training scheme temporally causal.
The second stage is an interface continuity loss defined on a dedicated sampling set near the window boundary. This term compares the $E_{z}$, $H_{x}$, and $H_{y}$ fields produced by the current window against the frozen outputs of the previous window at interface-centered points $(x_{\text{int}},y_{\text{int}},t_{\text{int}})$, and applies an additional penalty to their discrepancy. This enforces continuity more strongly over a narrow spatio-temporal band around the interface, rather than relying solely on a single IC slice defined at the window start. The weights $W_{\text{IC}}$ and $W_{\text{INTERFACE}}$ are used to tune the relative influence of these two continuity layers within the overall loss.
When applied together, these two continuity losses markedly suppress the artificial energy jumps expected at window transitions. In the window continuity analysis, the fact that the jumps in normalized total energy remain at an average level of $0.0028\%$ and at $0.0077\%$ in the worst case indicates that, despite the time-marching structure, the global solution can be stitched smoothly in both field and energy. This design plays a decisive role in reducing, in practice, the cumulative energy drift and window-transition discontinuities commonly observed in time-marched PINN architectures.
3.2.2 pinnL
To test whether the stable field and energy behavior exhibited by pinnA in the lossless PEC cavity is preserved under lossy medium conditions, a lossy variant pinnL was trained using the same time-marching architecture and the same configuration.
The differences between pinnA and pinnL are not limited to the medium model; the lossy physics directly affects both the PDE residual and the Poynting constraint formulations. The conductivity parameter is set to $\eta=\sigma/\varepsilon_{0}=0.5$ in non-dimensional form. The conductivity term $\eta E_{z}$ is incorporated into the PDE residual derived from Ampere’s law:
@@
A difference also exists in the loss function configuration. While the Poynting loss weight is $w_{\text{pyt}}=50$ in pinnA, this value is set to $w_{\text{pyt}}=15$ in pinnL. The rationale for this reduction is that in the lossy scenario, the Poynting residual inherently converges toward a nonzero target (the dissipation rate), and consequently, a constraint as aggressive as in the lossless scenario would hinder optimization.
Table 4: pinnL: Lossy Scenario Results
Time NRMSE L2 FDTD.E pinnL.E Rel.Enrgy.Err
@@
In terms of field accuracy, pinnL performs at the same order of magnitude as pinnA. The average $\text{NRMSE}_{\text{Total}}$ is measured as 0.08% in both scenarios, and the average NRMSE values for the $E_{z}$, $H_{x}$, and $H_{y}$ components remain on the order of one per mille throughout the entire time interval. The relative $L^{2}$ error averages 1.11% in the lossy scenario, remaining close to the lossless result. These results indicate that, despite the addition of the conductivity term and the reduction in the Poynting weight, the network retains its capacity to produce a consistent field solution across the spatio-temporal domain.
The energy metrics provide a more critical physical validation. The initial energy nearly coincides with FDTD in both scenarios, with the relative initial mismatch not exceeding 0.002%. As time progresses, the PINN and FDTD energies drift together in a consistent manner. The success criterion here is not that the global energy remains constant, but that the expected dissipation curve under the defined lossy dynamics is accurately tracked. The energy decreases from 0.0157 to 0.0059, corresponding to an approximately 63% reduction by $t=2.0$, and the PINN tracks this decay profile with a relative error below 0.2
Window transitions also remain controlled in the lossy scenario. While pinnA exhibits an average energy jump of 0.0028% and a maximum jump of 0.0077%, the corresponding values in pinnL are 0.0027% and 0.0095%, respectively. These results support the conclusion that the interface continuity mechanism preserves inter-window coherence in the lossy regime as well, keeping time-marching-induced artificial energy jumps at negligible levels.
3.2.3 pinnW
@@
Table 5: pinnW: Four-Window Variant of pinnA
Time NRMSE L2 FDTD.E pinnW.E Rel.Enrgy.Err
@@ 3.2.4 The Parenthesis Effect
In this study, an additional control variant was observed that was not initially targeted, yet produced a noteworthy divergence in training dynamics. This variant will hereafter be referred to as pinnP.
pinnP was obtained by making an exact copy of the pinnA folder and retraining it with the same training pipeline, the same random seed settings, and the same hyperparameters. The only difference between the two models is a single line within the Poynting loss function in the ‘losses.py‘ module. Apart from this, all files and all lines are identical character by character.
This difference can be summarized by the following two lines. In the pinnA configuration, the divergence term is written with parentheses grouping two subexpressions:
@@
In the pinnP configuration, the same summation is written without parentheses:
@@
Algebraically, the two expressions are equivalent. Nevertheless, the post-training metrics exhibited a noticeable separation, particularly in energy behavior. This observation is referred to in this study as the parenthesis effect.
Under the 20-window training setup, both pinnA and pinnP do not exhibit a pronounced cumulative energy drift pattern. However, pinnP shows weaker performance than pinnA in energy conservation metrics; the normalized energy curve oscillated over a wider band, and energy consistency metrics such as relative mismatch reached higher values.
The divergence becomes more apparent in the 4-window training setup. Although the 4-window variant of pinnA was introduced as pinnW in the previous section, the name pinnA will be used for the 4-window variant in this section as well to reduce confusion, since the comparison here focuses on the parenthesis effect rather than the window configuration.
In this configuration, pinnA experienced a noticeable transient overshoot around $t\approx 0.5$, yet exhibited a recovery behavior that brought the energy back toward the reference curve in the later portion of the simulation. The final relative energy error of the model at $t=2.0$ is approximately 0.236%.
In contrast, pinnP did not exhibit the same type of overshoot around $t\approx 0.5$; instead, it produced a more pronounced energy drift pattern along the time axis, and its relative energy error remained at approximately 0.416% at $t=2.0$.
Figure 4 compares, for the four-window training setup, the normalized energy curves of the two variants and the evolution of their relative mismatch.
Figure 3: pinnA vs pinnP (4-window versions)
A plausible contributing factor is floating-point non-associativity and its interaction with the autodiff graph. However, within the context of this study, the point to be emphasized is not the technical cause of the parenthesis effect, but rather its practical consequences on PINN models.
In classical numerical methods such as FDTD or FEM, writing two algebraically equivalent expressions in different forms in code does not alter the result, aside from negligible differences at the level of rounding error. PINNs, however, possess a fundamentally different structure due to their foundation on gradient-based optimization. Every term in the loss function influences the model parameters through the computational graph constructed during backpropagation. The topology of this graph can vary depending on how expressions are written. Consequently, two formulations that are mathematically identical may exhibit different characteristics in terms of optimization dynamics and final model behavior.
This observation points to an important consideration for PINN implementations: code-level modifications that appear algebraically inconsequential can affect the long-term behavior of the model in unexpected ways. Therefore, the parenthesis effect reported in this section should be interpreted not as an ablation experiment, but rather as an implementation note regarding the distinctive sensitivity of PINN-based solvers that sets them apart from classical numerical methods.
@@ 3.3 Ablation
Figure 4: Ablation study: Normalized energy evolution for pinnA (local Poynting), pinnGP (global Poynting), and pinnWP (no Poynting constraint) compared against FDTD reference.
The purpose of this ablation study is to isolate the cumulative energy drift behavior observed in the time-windowed training approach while keeping all other components fixed. To this end, three variants were trained using the same PINN architecture, the same time-marching scheme, the same sampling strategy, and the same loss weights: the reference model pinnA, which employs a local Poynting-based energy constraint; pinnGP, which uses a global Poynting regularizer defined via Gauss-Legendre quadrature; and pinnWP, in which the energy-conservation term is removed entirely.
@@
The three variants differ exclusively in how the Poynting-based energy constraint $\mathcal{L}_{\text{pyt}}$ is formulated and integrated into the total loss function. All other loss components (PDE, BC, IC, interface continuity) remain identical. Since all other aspects of training are held constant, the differences reported in Table 6 can be attributed directly to this design choice.
@@
The accuracy metrics reported in Table 6 indicate that all three variants achieve comparable field accuracy against the FDTD reference. Along the time axis, both NRMSE and $L^{2}$ errors remain at the sub-percent level, with only minor fluctuations among pinnA, pinnGP, and pinnWP. This similarity is important because it shows that the differences observed in the energy metrics cannot be explained by the field solution being better or worse. Instead, the divergence in energy behavior is primarily a direct consequence of the energy-regularization design.
When the energy metrics are examined, pinnA exhibits the most consistent behavior with the FDTD solution in terms of total energy. In the lossless cavity configuration, pinnA keeps the normalized total energy within a narrow band throughout the simulation and achieves the lowest energy-conservation error, both on average and at extreme values. In contrast, the typical cumulative energy drift behavior associated with time-windowed training becomes pronounced in pinnWP, where no energy constraint is present: small energy losses that occur within each window accumulate over time, causing the normalized total energy to decrease monotonically and causing the predicted field amplitudes to gradually diminish.
The global Poynting regularizer used in pinnGP only partially slows this drift and does not suppress it. Table 6 shows that pinnGP attains the highest values in both energy mismatch and energy variation compared to pinnA and pinnWP. Consequently, among the three models that lie in a similar accuracy band, a clear energy hierarchy emerges: energy conservation is best for pinnA, intermediate for pinnWP, and worst for pinnGP.
@@
To test whether this finding is merely an artifact of the energy functional defined on the continuous collocated grid, the analysis is also carried out on the native Yee grid discretization of the FDTD solver. The PINN fields are sampled on the Yee grid, and discrete energy time series are computed using the same discrete energy functional as FDTD. The energy-mismatch statistics derived from this analysis yield a ranking consistent with the continuous-grid results: pinnA produces the lowest mean and maximum differences as well as the lowest energy variation, these discrepancies increase for pinnWP and reach their highest level for pinnGP. Therefore, pinnA most accurately reproduces not only the continuous energy integral but also the discrete energy definition used by FDTD.
@@
The underlying reason for this behavior lies in the level and form at which Poynting’s theorem is integrated into the loss function.
pinnA: Local Poynting Residual.
The differential form of Poynting’s theorem is enforced as an additional conservation equation at the same collocation points as the Maxwell residuals. As a result, the condition $\partial_{t}u+\nabla\cdot\mathbf{S}=0$ is penalized directly at every point in the spatio-temporal domain rather than only at the integral level. Since the relationship between energy flux and energy density is locally constrained, each time window can establish its own internal energy balance, and the drift that would otherwise accumulate across window boundaries is largely suppressed. Consequently, the normalized energy curve exhibits a behavior close to the nearly flat profile expected for a lossless cavity, closely tracking the mildly oscillatory energy evolution of the FDTD solution.
pinnGP: Global Poynting Regularizer.
Instead of enforcing energy conservation pointwise, pinnGP constrains the rate of change of total energy integrated over the spatial domain. For a lossless PEC cavity, the boundary flux vanishes ($\mathbf{S}\cdot\hat{n}=0$ since $E_{z}=0$ on all walls), and Poynting’s theorem reduces to
$\frac{d}{dt}\int_{\Omega}u\,dA=0$ (25)
This integral is approximated using Gauss-Legendre quadrature with $N_{q}$ nodes and weights $\{(x_{j},y_{j}),w_{j}\}_{j=1}^{N_{q}}$:
The key distinction is that this formulation produces a single scalar constraint per time instant rather than $N_{c}$ independent constraints distributed across the spatio-temporal domain.
In such a setting, excess energy in one region of the domain can partially compensate for an energy deficit elsewhere. While optimization can keep the global integral within an acceptable range through cancellation of errors, local energy densities may still develop physically inconsistent distributions. In the context of time-windowed training, this mechanism makes it harder to suppress cumulative energy drift, because it allows local errors within each window to spread into the global integral rather than being corrected where they occur.
pinnWP: No Poynting Constraint.
The energy-conservation term is omitted entirely:
$\mathcal{L}_{\text{pyt}}=0.$ (27)
This variant serves as a control to isolate the effect of energy regularization from other components of the training pipeline.
This ablation study demonstrates that the cumulative energy drift observed in time-windowed PINN training is an independent failure mode that is not directly tied to field accuracy. It also quantitatively shows that physics-based energy constraints must be designed with care, not only in terms of whether they are included, but also in terms of the local versus global formulation and the manner in which they are integrated into the loss function. In particular, the pinnA configuration with a local Poynting regularizer provides a practically deployable solution for obtaining an energy-stable time-marching PINN formulation in this setting.

### v2

@@ 3.2.1 Main Model
pinnA is a fully-connected neural network trained along the time axis using a time marching approach. The network architecture consists of 8 hidden layers with 128 neurons per layer. The computational domain is defined by $x\in[0,1]$ and $y\in[0,\beta]$, where $\beta=1.0$ represents the aspect ratio of the cavity. The global time interval $t\in[0,2.0]$ is partitioned into 20 consecutive windows, each with width $\Delta t_{\text{window}}=0.10$, and trained independently. The initial condition is specified by a Gaussian pulse with amplitude $A=1.0$ and width $\sigma=0.1$. For each window, the training data consists of $100000/\beta$ collocation points, $20000/\beta$ boundary points, and $3000/\beta$ initial condition points. The loss function is defined as a weighted sum of five terms: PDE residual ($w_{\text{PDE}}=1.0$), boundary conditions ($w_{\text{BC}}=50.0/\beta$), initial condition ($w_{\text{IC}}=100.0$), window-interface continuity ($w_{\text{interface}}=75.0$), and Poynting-vector-based energy conservation ($w_{\text{Poynting}}=10.0$). Each window is trained using 1500 epochs of the Adam optimizer with learning rate $10^{-3}$, followed by 300 epochs of L-BFGS with learning rate $0.5$.
@@
Field-level accuracy of pinnA relative to FDTD is assessed using the Normalized Root Mean Square Error (NRMSE) and the relative $L^{2}$ norms defined in the previous section. The results for selected time steps are reported in the Table 1. The reported average values are computed not only over the time steps shown in the Table 1, but as an average over all time steps used in the analysis, $t=[0.00,0.05,0.10,0.15,\ldots,2.00]$.
As shown in the Table 1, the NRMSE values for all three field components remain on the order of 0.1–0.2% throughout the full time interval. Even the relative $L^{2}$ error reaches only about $1.33\%$ in the worst case and stays around $1.01\%$ on average.
@@
Since the values in the Table 2 and the figure are computed on a collocated grid for the PINN solution, FDTD and PINN are additionally evaluated on a common Yee grid to confirm consistency. Under the staggered-grid layout, the mean absolute difference over time is $3.74\times 10^{-6}$, and the maximum difference is $9.07\times 10^{-6}$. When normalized by the reference energy, the mean relative mismatch is only $0.024\%$, and the maximum mismatch is $0.057\%$. In other words, the discrete energy functional of the PINN tracks the discrete FDTD energy across the entire spatio-temporal domain at a level well below $10^{-3}$ relative error.
@@
3.2.2 Lossy Scenario
To test whether the stable field and energy behavior exhibited by pinnA in the lossless PEC cavity is preserved under lossy medium conditions, a lossy variant was trained using the same time-marching architecture and the same configuration.
The differences between pinnA and the lossy variant are not limited to the medium model; the lossy physics directly affects both the PDE residual and the Poynting constraint formulations. The conductivity parameter is set to $\eta=\sigma_{c}/\varepsilon_{0}=0.5$ in non-dimensional form. The conductivity term $\eta E_{z}$ is incorporated into the PDE residual derived from Ampere’s law:
@@ 3.2.2 Lossy Scenario
A difference also exists in the loss function configuration. While the Poynting loss weight is $w_{\text{pyt}}=10$ in pinnA, this value is set to $w_{\text{pyt}}=15$ in the lossy variant.
Table 4: Lossy Scenario Results
Time NRMSE L2 FDTD.E pinn.E Rel.Enrgy.Err
@@
In terms of field accuracy, the lossy variant performs at the same order of magnitude as pinnA. The average $\text{NRMSE}_{\text{Total}}$ is measured as 0.08% in both scenarios, and the average NRMSE values for the $E_{z}$, $H_{x}$, and $H_{y}$ components remain on the order of one per mille throughout the entire time interval. The relative $L^{2}$ error averages 1.11% in the lossy scenario, remaining close to the lossless result. These results indicate that, despite the addition of the conductivity term and the reduction in the Poynting weight, the network retains its capacity to produce a consistent field solution across the spatio-temporal domain.
The energy metrics provide a more critical physical validation. The initial energy nearly coincides with FDTD in both scenarios, with the relative initial mismatch not exceeding 0.002%. As time progresses, the PINN and FDTD energies drift together in a consistent manner. The success criterion here is not that the global energy remains constant, but that the expected dissipation curve under the defined lossy dynamics is accurately tracked. The energy decreases from 0.0157 to 0.0059, corresponding to an approximately 63% reduction by $t=2.0$, and the PINN tracks this decay profile with a relative error below 0.2%.
Window transitions also remain controlled in the lossy scenario. While pinnA exhibits an average energy jump of 0.0028% and a maximum jump of 0.0077%, the corresponding values in the lossy variant are 0.0027% and 0.0095%, respectively. These results support the conclusion that the interface continuity mechanism preserves inter-window coherence in the lossy regime as well, keeping time-marching-induced artificial energy jumps at negligible levels.
3.2.3 Four-Window Variant
@@
Table 5: Four-Window Variant of pinnA
Time NRMSE L2 FDTD.E pinn.E Rel.Enrgy.Err
@@
3.2.4 High Frequency Scenario
To test the robustness of pinnA against spectral bias, tests were also conducted in the high frequency region. In the pinnA configuration, the Gaussian pulse width $\sigma$ = 0.1 was set, which corresponds to a wave structure of approximately 7.8 periods in the medium frequency regime. In the high frequency variant pinnHF, $\sigma$ = 0.05 was set, thereby increasing the maximum dominant frequency from 3.902 to 8.780 and the number of periods to approximately 17.6.
Table 6: Comparison of pinnA and pinnHF configurations.
Metric pinnA pinnHF
NRMSE (avg) 0.09% 0.10%
$L^{2}$ Error (avg) 1.01% 2.21%
Energy Rel. Error 0.02% 0.21%
Energy Mismatch (Yee Grid) 0.02% 0.21%
Max. Dominant Freq. 3.902 8.780
Max. Number of Periods $\sim$7.8 $\sim$17.6
When the results are examined, pinnHF demonstrated strong performance despite the frequency content increasing by more than twofold. The average NRMSE was measured as 0.10%, the average $L^{2}$ error as 2.21%, and the relative energy error as 0.21%. These results indicate that the proposed methodology works effectively in the high frequency regime as well, despite the spectral bias phenomenon where neural networks tend to learn low frequency components faster. The fact that error metrics remain within acceptable limits despite the significant increase in frequency content suggests that the time marching strategy and Poynting based regularization can partially compensate for spectral bias.
When the frequency band is increased further, it was observed that the model’s success diminishes. For $\sigma$ = 0.03, when the maximum number of periods is approximately 29.9, the average relative energy error reaches 6.862%. A large portion of the relative error occurs at the initial instant when the Poynting regularizer operates with low weight. With the active engagement of the Poynting regularizer, a decrease in energy drift was observed.
It can be stated that $\sigma$ = 0.05 represents a performance boundary for pinnA, beyond which the current model loses its stability. However, it has been understood that the spectral bias problem can be substantially mitigated through improvements to the Poynting regularizer. A spectral bias resistant PINN configuration is targeted in future work.
@@ 3.2.5 The Parenthesis Effect
In this study, an additional control variant was observed that produced a noteworthy divergence in training dynamics. This variant will hereafter be referred to as pinnP.
pinnP was obtained by copying the pinnA folder exactly and retraining with the same model. The only difference is a single line in the Poynting loss function. Apart from this, all files are identical character by character.
This difference can be summarized by the following two lines. In the pinnA, the divergence term is written with parentheses grouping two subexpressions:
@@
In the pinnP, the same line is written without parentheses:
@@
Algebraically, the two expressions are equivalent. Nevertheless, post-training metrics exhibited a noticeable separation, particularly in energy behavior. This observation is referred to as the parenthesis effect.
Under the 20-window setup, both pinnA and pinnP avoid pronounced cumulative energy drift. However, pinnP shows weaker energy conservation; the normalized energy curve oscillated over a wider band and relative mismatch reached higher values.
The divergence becomes more apparent in the 4-window setup. Here, pinnA experienced a transient overshoot around $t\approx 0.5$, yet recovered toward the reference curve later. Its final relative energy error at $t=2.0$ is approximately 0.236%.
In contrast, pinnP did not exhibit such overshoot; instead, it produced a more pronounced drift pattern, with relative energy error remaining at approximately 0.416% att $t=2.0$.
Figure 4 compares the normalized energy curves and relative mismatch evolution for the four-window setup.
Figure 3: pinnA vs pinnP (four-window versions)
A plausible contributing factor is floating-point non-associativity and its interaction with the autodiff graph. However, the point here is not the technical cause but the practical consequences for PINN models.
In classical methods such as FDTD or FEM, writing algebraically equivalent expressions differently produces negligible differences at rounding-error level. PINNs, however, are built on gradient-based optimization, where every loss term influences parameters through the computational graph. The graph topology can vary with how expressions are written, so mathematically identical formulations may exhibit different optimization dynamics.
This observation highlights an important consideration: code-level modifications that appear algebraically inconsequential can affect long-term model behavior unexpectedly. The parenthesis effect should therefore be interpreted not as an ablation experiment, but as an implementation note on the distinctive sensitivity of PINN-based solvers compared to classical methods.
@@ 3.3 Ablation
Figure 4: Ablation: pinnA (local Poynting), pinnGP (global Poynting), and pinnWP (no Poynting constraint) compared against FDTD.
@@
The purpose of this ablation study is to isolate the cumulative energy drift behavior observed in the time-windowed training approach while keeping all other components fixed. To this end, three variants were trained using the same PINN architecture, time-marching scheme, sampling strategy, and loss weights. The variants differ exclusively in how the Poynting-based energy constraint $\mathcal{L}_{\text{pyt}}$ is formulated; all other loss components (PDE, BC, IC, interface continuity) remain identical. Since all aspects of training other than this single design choice are held constant, the differences reported in Table 7 can be attributed directly to the energy-regularization strategy.
pinnA: Local Poynting Residual.
The differential form of Poynting’s theorem is enforced as an additional conservation equation at the same collocation points as the Maxwell residuals. The condition $\partial_{t}u+\nabla\cdot\mathbf{S}=0$ is penalized at every point in the spatio-temporal domain. Since the relationship between energy flux and energy density is locally constrained, each time window can establish its own internal energy balance, and the drift that would otherwise accumulate across window boundaries is largely suppressed.
pinnGP: Global Poynting Regularizer.
Instead of pointwise enforcement, pinnGP constrains the rate of change of total energy integrated over the domain, approximated via Gauss–Legendre quadrature with $N_{q}$ nodes and weights $\{(x_{j},y_{j}),w_{j}\}_{j=1}^{N_{q}}$:
This formulation produces a single scalar constraint per time instant rather than $N_{c}$ independent constraints distributed across the domain. Consequently, excess energy in one region can compensate for a deficit elsewhere at the integral level, masking local violations through cancellation of errors. As discussed below, this mechanism does not merely limit the effectiveness of the constraint; it can actively degrade energy behavior relative to having no constraint at all.
pinnWP: No Poynting Constraint.
The energy-conservation term is omitted entirely ($\mathcal{L}_{\text{pyt}}=0$). This variant serves as a control to isolate the effect of energy regularization from other components of the training pipeline.
@@
The accuracy metrics in Table 7 indicate that all three variants achieve comparable field accuracy against the FDTD reference. Both NRMSE and $L^{2}$ errors remain at the sub-percent level along the time axis, with only minor fluctuations among the three models. This similarity is important: the differences in energy behavior cannot be attributed to one model producing better or worse field solutions, but are a direct consequence of the energy-regularization design.
When the energy metrics are examined, a clear hierarchy emerges. pinnA exhibits the most consistent energy behavior, keeping the normalized total energy within a narrow band and achieving the lowest conservation error both on average and at extreme values. pinnWP, despite having no energy constraint, occupies an intermediate position: cumulative drift is visible as the normalized energy decreases monotonically, but the magnitude remains moderate. The counterintuitive result is that pinnGP, the only variant with a global energy constraint, performs worst in every energy metric. Its average relative error (0.088%) is roughly double that of pinnWP (0.044%) and nearly four times that of pinnA (0.024%). The Cons.PINN column makes the contrast even more explicit: by $t=2.0$, pinnGP has drifted to $0.142\%$, compared with $0.095\%$ for pinnWP and only $0.031\%$ for pinnA.
This outcome can be understood through the cancellation-of-errors mechanism inherent to the global formulation. Because the global loss penalizes only the net integral $\sum w_{j}\partial_{t}u$, the optimizer can satisfy the constraint by allowing positive and negative $\partial_{t}u$ violations to cancel across the domain. The resulting locally inconsistent energy distributions do not trigger a penalty at the loss level, yet they distort the solution in ways that compound across window boundaries. In effect, the global constraint introduces an additional optimization pathway that permits, and may even encourage, physically inconsistent local energy distributions, producing worse cumulative drift than the unconstrained baseline.
@@
To verify that this finding is not an artifact of the collocated evaluation grid, the analysis is repeated on the native Yee grid discretization of the FDTD solver. The PINN fields are sampled at Yee grid locations and discrete energy time series are computed using the same discrete energy functional as FDTD. The resulting ranking is fully consistent: pinnA produces the lowest mean and maximum differences as well as the lowest energy variation; these discrepancies increase for pinnWP and reach their highest level for pinnGP.
@@
This ablation study demonstrates two key findings. First, cumulative energy drift in time-windowed PINN training is an independent failure mode not directly tied to field accuracy. Second, physics-based energy constraints must be designed with care: the global Poynting formulation not only fails to suppress drift but actively worsens it relative to having no constraint, whereas the local formulation used in pinnA provides effective energy stabilization. The critical distinction lies in whether the constraint operates at the pointwise or integral level, a design choice that determines whether local energy violations are corrected where they occur or allowed to spread and compound through the cancellation-of-errors mechanism.